OpenStudy (goformit100):

A boy was asked to find the LCM of 3,5,12 and some another number. But while calculating he wrote 21 instead of 12 and yet got the correct answer. What would be the forth number.

OpenStudy (goformit100):

@AcidRa1n

OpenStudy (zzr0ck3r):

5 does not divide 12 so 12 is not the least common multiple of 5 and any number...

OpenStudy (goformit100):

okay

OpenStudy (goformit100):

Don't Spam My Question^

OpenStudy (anonymous):

sorry didn't realize one message was spam.. But ok

OpenStudy (goformit100):

@zzr0ck3r After calculation I got 28 is it correct ? it was a Long Paper working.

OpenStudy (zzr0ck3r):

this question does not make sense to me lcm mean least common multiple that means that the lcm of the set {a,b,c,d} will be divisible by each number in the set so the lcm of 3,5,12,some other number must be divisible by each of these numbers 5 does not divide 12 so 12 cant be the lcm of this set....

OpenStudy (goformit100):

Okay. Then what to do then ?

OpenStudy (zzr0ck3r):

no idea, the question does not make sense \[lcm\{3,5,12\} = lcm\{3,5,3*2^2\}=3*5*4=60\]

OpenStudy (zzr0ck3r):

so no matter what number we add to the set, this is only going to get bigger or stay the same......